OpenStudy (anonymous):

The polynomial 8x2 - 9 is a difference of squares. T/F?

OpenStudy (anonymous):

8x^2-9**

OpenStudy (anonymous):

I think true but people are saying false?

OpenStudy (whpalmer4):

Difference of squares has the form \(a^2-b^2\) Can you find a value for \(a\) and \(b\) to match your equation? Obviously, \(b=3\) is half of the job, what about for \(a^2 = 8x^2\)?

OpenStudy (anonymous):

?? All I asked was "True or False"

OpenStudy (whpalmer4):

Yes, and I'm trying to help you know the answer without asking us :-)

OpenStudy (whpalmer4):

\[a^2=8x^2\]Can you solve that for \(a\)?

OpenStudy (anonymous):

I'm going with false.

OpenStudy (whpalmer4):

Take the square root of both sides: \[\sqrt{a^2} = \sqrt{8x^2}\]Remember that \(\sqrt{a^2} = a\) \[a = \sqrt{8}*\sqrt{x^2} = x\sqrt{8}\] So we can see that it is TRUE, because \[(x\sqrt{8})^2-3^2 = 8x^2-9\]